### v1

# Fragile Preferences: A Deep Dive into Order Effects in Large Language Models

Abstract
Large language models (LLMs) are increasingly used in decision-support systems across high-stakes domains such as hiring and university admissions, where decisions often involve selecting among competing alternatives. While prior work has noted positional order biases in LLM-driven comparisons, these biases have not been systematically dissected or linked to underlying preference structures. We provide the first comprehensive investigation of positional biases across multiple LLM architectures and domains, uncovering strong and consistent order effects—including a novel centrality bias not previously documented in human or machine decision-making. We also find a quality-dependent shift: when options are high quality, models exhibit primacy bias, but favor latter options when option quality is low. We further identify a previously undocumented bias favoring certain names over others. To distinguish superficial tie-breaking from true distortions of judgment, we introduce a framework that classifies pairwise preferences as robust, fragile, or indifferent. We show that order effects can lead models to select strictly inferior options, and that positional biases are typically stronger than gender biases. These findings suggest that LLMs are not merely inheriting human-like biases, but exhibit distinct failure modes not seen in human decision-making. We propose targeted mitigation strategies, including a novel use of the temperature parameter, to reduce order-driven distortions.
1 Introduction
Large language models (LLMs) are increasingly integrated into decision-support systems across high-stakes domains such as hiring, healthcare, and loan approvals [1, 2, 3]. In these contexts, ensuring fairness and transparency is not just an ethical imperative but often a legal requirement, as recent regulations emphasize algorithmic accountability and nondiscrimination [4, 5, 6]. Among these domains, hiring has received particular attention due to both its societal importance and the real-world use of LLMs in resume screening [7, 6].
To date, most studies of LLMs in hiring have focused on scoring individual resumes or responses, documenting biases related to gender, race, and socioeconomic background [8, 7, 9, 10, 11, 12, 13]. A growing line of work now advocates for pairwise or setwise comparisons, arguing that relative judgments are more consistent and more human-like [14, 15, 16, 17]. However, as these frameworks gain traction, it becomes increasingly urgent to examine the biases they may introduce, especially in high-stakes settings where even small distortions can have large consequences.
Ranking-based decisions are particularly susceptible to positional and contextual effects. In human judgment, these include primacy and recency [18, 19, 20, 21, 22, 23], anchoring [24, 25], contrast effects [26, 27], and decoy effects [28, 29, 30]. These biases can undermine consistency and lead to inconsistent or irrational preferences. Recent studies show that LLMs are vulnerable to similar biases. Positional effects have been documented in multiple-choice settings [31, 32], classification tasks [33, 34], and pairwise evaluations [17, 35]. Other context-related effects, such as anchoring and decoy effects, have also been observed [36, 37]. Yet this literature remains largely descriptive: we know these effects exist, but not how they manifest in different contexts, how they interact with other biases, or what mechanisms drive them.
In this work, we systematically investigate positional biases in LLMs. We uncover previously unreported biases, show that these effects can meaningfully distort underlying preferences, and find important differences from human biases. Across two domains—hiring and color selection—we uncover a consistent, quality-dependent pattern in pairwise comparisons: when the options are high quality, LLMs exhibit a primacy bias, favoring the first candidate or color. For lower-quality options, however, they favor latter options. In triplewise comparisons, we identify a novel centrality bias: a consistent preference for the middle option. Both the centrality effect and the quality-dependent shift from primacy to recency appear to be unique to LLMs and have not, to our knowledge, been reported in human (or LLM) decision-making. We further show that positional biases are, for the most part, stronger than gender biases, and uncover a strong and previously undocumented source of distortion: a bias favoring certain names over others. These patterns suggest that LLMs are not merely inheriting human-like heuristics from their training data, but are manifesting new failure modes with distinct underlying mechanisms.
Separate from these novel patterns, we highlight a conceptual distinction that has been largely overlooked in prior work on order effects in both humans and LLMs. Positional biases are typically treated as a single phenomenon, but they can arise from two fundamentally different mechanisms: tie-breaking heuristics and true preference distortions. Tie-breaking occurs when the agent (human or AI) has no clear preference and defaults to a fixed positional rule (e.g., always guessing option C on a multiple-choice test). In contrast, a preference distortion occurs when the order of presentation alters an otherwise strict preference. This distinction matters: while tie-breaking is relatively harmless, preference distortion implies the model may select inferior options due solely to presentation order.
Disentangling these mechanisms is difficult, especially in domains where ground-truth preferences are ambiguous. Many prior studies suggest that order effects emerge mainly when options are similar [31, 17, 32], consistent with the tie-breaking view. However, we show that positional effects in LLMs can lead to genuine preference reversals. Leveraging the ability to repeatedly query LLMs and control their sampling behavior via temperature, we introduce a simple yet effective mitigation strategy: querying the model several times at higher temperature settings. When the temperature parameter is increased, the models’ output becomes more stochastic, and repeated sampling shows more variation. We can quantify the model’s preferences by the frequency with which the options are selected: when the probability of selecting option $a$ over $b$ at temperature 1 is statistically significant, we infer that the model prefers $a$ to $b$, even if the model appears indifferent at temperature 0. In such scenarios, increasing the temperature can improve accuracy, contrary to the prevailing understanding that greater accuracy is attained at lower temperature settings[38, 39]. We elaborate on this technique in Subsection 2.3.
To formalize the distinction between tie-breaking and true distortions, we extend the classical framework of decision-making, which represents preferences using the binary relations $a\succ b$ (strict preference) and $a\sim b$ (indifference). We introduce a refined notation that distinguishes between two types of strict preferences. We write $a\ggcurly b$ to indicate that the preference for $a$ over $b$ is strong and stable across contexts. In contrast, $a\succcurlyeq b$ denotes a genuine but unstable preference that may be reversed by order effects. This framework allows for clearer interpretation of observed behavior. if $a\succcurlyeq b$, but $b$ is consistently selected when presented first, we attribute the reversal to a distortion of the underlying preference, not to heuristic tie-breaking. However, if $a\sim b$, the cause for the order effect cannot be disambiguated.
Together, our findings suggest that LLM comparisons are governed by both familiar and novel biases, and that these biases can systematically skew decisions. The implications are broad: positional distortions often outweigh demographic biases, and they cannot be reliably corrected by averaging or randomizing presentation. In our discussion, we explore both theoretical explanations and practical mitigation strategies.
2 Results
(a) Proportion of color choices by position across 4 quality tiers in pairwise comparisons. Tier quality is decreasing from left to right: ‘Plain’ is the highest quality tier and ‘Harsh’ is the lowest.
(b) Positional bias in pairwise resume selection, aggregated across professions.
Figure 1: Positional bias in pairwise comparisons at $T=1$. (a) shows the effect of presentation order on color selection in pairwise comparisons for each quality tier. (b) shows aggregate positional effects in resume evaluations across professions. In both tasks, higher-quality options tend to exhibit a primacy effect and lower-quality ones a recency effect, though the precise threshold separating them varies by model and domain. Full results for color comparisons are provided in Appendix C ; results for resumes disaggregated by profession are given in Appendix D
We evaluated positional biases in three widely used large language models—GPT-4o-mini, Claude 3 Haiku, and Llama 3 8B—across two domains designed to reflect real-world comparative decisions: resume screening and color selection. As order effects are most likely to arise when alternatives are similar in quality [17, 31], we constructed sets of items that are close in overall quality and grouped them into four tiers: ‘Ideal’, ‘Fair’, ‘Plain’ and ‘Harsh’ for colors and ‘Best’, ‘Good’, ‘Mediocre’ and ‘Weak’ for resumes. In the hiring domain, we used LLM-generated resumes spanning the four quality tiers in four professions. In the color domain, we curated sets of wall paint colors grouped by their suitability for a child’s room.
For each model and domain, we conducted exhaustive pairwise and triplewise comparisons, testing all permutations of item order. Prompts included “Select the strongest candidate” and “Which color is best for a kid’s room?” Each model was evaluated at multiple temperature settings. At temperature 0 ($T=0$), models behave almost deterministically, [^1] While temperature 0 is often assumed to produce deterministic outputs, we observed occasional variability in all models even at $T=0$. selecting the most likely answer. At temperature 1 ($T=1$), responses vary across repeated queries, enabling statistical analysis of positional effects. In our analyses, we treated each $T=1$ response as an independent sample from the model’s output distribution. Full experimental details appear in Section 4 (Materials and Methods) and the Appendix.
2.1 Order Effects in Pairwise Comparisons
We queried GPT-4o-mini, Claude 3 Haiku, and Llama 3 8B on exhaustive pairwise comparisons in two domains—resumes and colors—at multiple temperature settings. In every setting, higher-quality tiers exhibited a primacy effect (more likely to be selected when presented first), while lower-quality tiers showed a recency effect (more likely to be selected when presented second); these effects are clearly visible in Figure 1. For example, when comparing the two colors of the ‘Fair’ tier (Gentle Coral and Buttercream Yellow), GPT-4o-mini chose the first option 100% of the time at $T=0$ and 89% at $T=1$ (binomial $p=2.5\times 10^{-16}$). The quality tier at which bias flipped from primacy to recency was highest for Claude 3 Haiku and lowest for Llama 3 8B, a pattern consistent across both domains. Full statistics appear in Appendices C (colors) and D (resumes).
(a) Proportion of color choices by position across 4 quality tiers in triplewise comparisons.
(b) Aggregate positional bias in triplewise resume selection across 4 quality tiers.
Figure 2: Positional bias in triplewise comparisons at $T=0$. (a) shows the effect of presentation order on color selection. (b) shows aggregate positional effects in resume evaluations across professions. Additional results for color selection are provided in Appendix F, and resumes selection results disaggregated by profession are in Appendix G.
2.2 Order Effects in Triplewise Comparisons
In triplewise comparisons, higher-quality tiers exhibit a primacy effect, consistent with the pairwise results. For lower-quality tiers, all models favor latter positions, though they differ in whether the second or third option is preferred; see Figure 2. GPT-4o-mini and Claude 3 Haiku both display a centrality bias—favoring the middle option over first or last. For example, at $T=0$, GPT-4o-mini always selected the middle option for ‘Plain’ colors, and Claude 3 Haiku did so 83% of the time for ‘Harsh’ colors. In contrast, Llama 3 8B did not display a centrality bias, but instead showed simultaneous primacy and recency effects, resembling the serial position effects observed in humans [21, 40]: among colors in the ‘Fair’ tier at $T=0$, Llama 3 8B chose the first and last positions 50% of the time each, and never selected the middle option. See Appendices F (colors) and G (resumes) for additional details.
Across models and domains, the bias generally shifts monotonically toward later positions as quality declines, with the exception of Claude 3 Haiku in the color domain. Positional biases are notably weaker in resume selection for Claude 3 Haiku and Llama 3 8B compared to color selection. As we show in the Name Bias section below, this attenuation is largely due to interference from strong preferences for specific names.
2.3 Tie-Breaking Heuristics vs. True Distortions
[…]
(a) Name selection frequencies in triplewise comparisons. The histogram shows how often each candidate name was selected across all triplewise tasks: selection counts reflect the number of times the name was chosen as the best option. See Appendix H for additional results and details.
(b) Gender and order effects at $T=0$ in pairwise resume comparisons. All models show stronger biases for presentation order than for gender. Results by profession and temperature are provided in Appendix I .
Figure 3: Interaction of order effects with other biases. (a) shows the distribution of name selections in triplewise comparisons. Claude 3 Haiku exhibits strong preferences for certain names, while GPT-4o-mini shows a relatively balanced selection pattern. (b) shows the effect of gender and presentation order in pairwise resume comparisons. Order effects appear stronger than gender effects.
[…]
2.4 Interaction of Positional Bias with Other Biases
In addition to positional order effects, LLMs exhibit other well-documented biases, such as preferences based on gender or identity. In this section, we examine how these biases interact with positional effects. While gender bias in algorithmic decision-making is well documented, we also uncover an additional bias: consistent preferences for specific individual names, even when those names are carefully selected to avoid known bias-triggering cues such as race and ethnicity. To our knowledge, such name-specific biases have not been documented in human decision-making under comparable conditions. Across all models, we find that these biases—though model-specific in strength and direction—are generally weaker than the positional effects, although in some cases they appear to compete with or modulate them.
Name Bias.
The positional bias in triplewise color comparisons is so strong that in every tier, there is at least one position that is never chosen at $T=0$ (except in GPT-4o-mini’s ‘Fair’ tier, which selected the last position a single time). This positional skew appears less pronounced in resume comparisons—particularly for Claude 3 Haiku and Llama 3 8B. Upon closer inspection, this reduction in positional bias appears to be driven by strong, competing biases for or against specific synthetic personas, especially individual names. Claude 3 Haiku, in particular, exhibits strong name-based preferences: it consistently favors certain names over others, regardless of resume content or position. For example, in direct comparisons, Claude 3 Haiku ‘Christopher Taylor’ over ‘Andrew Harris’ in 64% of the cases (82 out of 128 trials, binomial $p=0.0019,h=0.29$). These patterns emerged even though all names were drawn from a controlled set of generic Caucasian U.S. identities, chosen to minimize racial, ethnic, and other socially marked biases that have been documented in LLMs [41, 42].
This pattern is evident in 3(a), which shows the distribution of selection frequencies for each name across triplewise comparisons. Claude 3 Haiku’s distribution is flat, indicating that a few names are chosen disproportionately often while others are rarely selected—consistent with strong and persistent name preferences. GPT-4o-mini, by contrast, displays a near-perfect normal distribution, suggesting minimal name-based bias in the triplewise setting. Llama 3 8B lies between these extremes, with moderate variation across names. Although GPT-4o-mini shows no signs of name bias in triplewise comparisons, we detect a small but statistically significant name preference in the pairwise setting. See Appendix H for full details.
Gender Bias.
Prior work on gender bias in algorithmic hiring is mixed: some studies report little to no bias [43], others find systematic preferences for male candidates [44, 7], while more recent evaluations suggest a tilt toward female candidates [8, 9] or context-dependent outcomes [13]. Studies using pairwise evaluations have found a bias favoring female candidates [35, 8, 9], a pattern we also replicate: for example, GPT-4o-mini selects female candidates 55.7% of the time at $T=1$ ($p=3.22\times 10^{-4}$, $h=0.11$), with Claude 3 Haiku and Llama 3 8B showing similar trends.
While gender bias in algorithmic decision-making is an important topic in its own right, our primary goal is to examine how gender interacts with positional order effects. We find that gender effects are overshadowed by substantially stronger order biases. Across all three models, the candidate in the dominant position (first or second, as defined by same-gender comparisons) is selected at much higher rates. For GPT-4o-mini, this rate reaches 92.8% in mixed-gender comparisons—compared to 93.6% in same-gender trials—suggesting that gender bias can at times counteract, but rarely override, the positional preference. Claude 3 Haiku and Llama 3 8B show similar patterns, with positional bias rates of 84.3% and 72.3% in mixed-gender comparisons, versus 87.6% and 89.2% in same-gender comparisons, respectively. Bootstrapped confidence intervals confirm that order bias significantly exceeds gender bias across all models (95% CI for $h_{\text{order}}-h_{\text{gender}}$ excludes 0 in all cases).
This difference between the effect sizes of the positional and gender effects can be clearly seen in 3(b), where gender effects manifest in two ways: a modest overall preference for female candidates, and the observation that nearly all instances of overcoming the positional disadvantage involve female candidates. Full statistics appear in Appendix I.
3 Discussion
This study demonstrates systematic positional biases in LLM-driven comparisons, with methodological, theoretical, and practical implications. Across both color and hiring domains, we find that positional biases are quality-dependent and strong: primacy effects dominate for high-quality options, while lower-quality choices elicit a recency effect.
Most prior research has not clearly distinguished between two possible sources of positional effects: arbitrary tie-breaking and genuine preference distortion. Many studies have observed that order effects appear when the model is indifferent between the different options [31, 17, 32], and studies on positional biases and the decoy effect often purposefully use alternatives of similar quality  [45, 37]. While a few human studies have attempted to disentangle these explanations [46], conclusions remain limited due to humans’ limited ability to determine their ‘true’ preferences [24] or compute expected utilities [47]. The difficulty to determine ground-truth preferences in human subjects is further complicated by the fact that preferences are influenced by contextual factors and shift over time [48].
By contrast, LLMs allow repeated querying under controlled conditions. We exploit this property, along with the temperature parameter (which has no human analogue), to uncover latent preferences and distinguish tie-breaking from distortion by querying across multiple temperatures. Using this methodology, we show that LLMs do not merely exhibit positional effects when indifferent but can reverse strict preferences based solely on presentation order. Although subtle, this distinction matters: tie-breaking is relatively benign, whereas distortion leads to selecting an inferior option purely because of the order of presentation.
Much of the work on LLM bias has centered on uncovering and correcting cognitive distortions, such as anchoring, confirmation bias, or gender stereotyping, that models acquire from human-generated training data [49, 50, 51, 41]. These studies often argue that models inherit cognitive biases that are present in the data [52, 53, 54]. Even studies that identify behavior that diverges from human norms (for instance, a female‐candidate preference in resume‐ranking tasks, where humans often favor male candidates) still frame these findings through the lens of human cognitive biases [8, 9]. In contrast, we uncover two new failure modes that have no clear analogue in human decision-making: a centrality bias in triplewise comparisons and a name bias, even though names were explicitly chosen to avoid known demographic triggers.
These patterns raise new theoretical questions. Classical explanations for human order effects, such as memory limitations [55, 56], proactive inhibition [57], or heuristics [24], cannot readily explain our findings of quality-dependent bias shifts or the centrality bias. One possible account—drawing on the long-established notions of reference points [40, 58] and contrast effect [26, 27] in behavioral decision theory—is that LLMs establish an implicit reference point from their training data and then evaluate each option in contrast to this baseline. Above-baseline options are inflated, leading to primacy, while below-baseline ones are penalized. However, it is not clear how to extend this explanation to biases observed in triplewise comparisons, in particular the centrality bias. Moreover, model-specific differences (e.g., centrality effects in GPT-4o-mini and Claude 3 Haiku but not Llama 3 8B), indicate that different architectures details may give rise to distinct bias mechanisms. Together, these findings call for new theoretical frameworks to explain how biases emerge and propagate in artificial agents—frameworks that go beyond human cognitive models and the biases embedded in training data.
From a practical standpoint, standard mitigation strategies such as randomizing orders or averaging across queries do not fully eliminate positional distortions. Instead, we propose approaches that leverage LLM-specific characteristics. First, raising the temperature uncovers fragile preferences that deterministic (zero-temperature) settings suppress. Although low temperatures are often favored for precision and factuality, and high temperatures for creativity [38, 39], we find that in these comparison tasks a higher temperature can actually improve selection accuracy. Second, positional bias itself can be informative: a pronounced primacy effect may indicate that both options are of high quality, whereas a strong recency effect may signal that both are of lower quality.
While our findings hold across three contemporary models, new architectures should be audited for emergent order and identity biases. In high-stakes domains such as legal reasoning, medical diagnosis, or scientific peer review, bias-aware evaluation pipelines can combine model-driven diagnostics with human-in-the-loop checks. Finally, causal interpretability methods and adversarial probing could illuminate the internal mechanisms that give rise to preference distortions, paving the way toward bias-robust LLM design and deployment.
4 Materials and Methods
4.1 Resume and Color Generation and Selection
As order effects are more pronounced when the options are of comparable quality, we generated three options for each of four quality tiers of colors and resumes. For colors, we denoted these tiers by ‘Ideal’, ‘Fair’, ‘Plain’ and ‘Harsh’; for resumes, by ‘Best’, ‘Good’, ‘Mediocre’ and ‘Weak’. We chose different names for the tiers in each domain so as not to imply an equivalence between them (i.e., ‘Plain’ colors are not necessarily comparable to ‘Mediocre’ resumes in terms of quality).
We first generated the candidates for each tier. For the resumes, we had GPT-4o-mini generate nine sets of four resumes of decreasing quality for each of four professions: Mechanical Engineer, Registered Nurse, Journalist, and Real Estate Agent. We removed the personal information from the resumes. The resume generation prompts are shown in Appendix B. For colors, we asked each model to generate colors that they ‘thought’ were comparable.
For each quality tier, we performed triplewise comparisons among the candidate options and selected three options based on the following criteria. First, each had to be selected in at least one of the six possible permutations. Second, to ensure that the tier labels were meaningful, we checked that colors from a higher tier were chosen over those from a lower tier in at least two-thirds of all pairwise comparisons. To select the colors for pairwise comparisons in each tier, we evaluated all three possible pairs among the three options and selected a pair in which neither color was consistently preferred over the other. We note that our selection methodology led to the fact that we did not necessarily use the same resumes or colors for the three models.
For the personal information, We generated 30 fictional personas, using 15 common American Caucasian last names, male first names and female first names, selected from the list of the most popular names in the US as tracked by the SSA [59]. We paired each first name with a random surname and state, and randomly extrapolated to a complete profile. For more details, see Appendix B.
4.2 Pairwise Comparisons
For each pair, we presented the model with both orders to isolate position effects. For resumes, we selected four male names and four female names and performed exhaustive same-gender comparisons. That is, for each pair of names $i$ and $j$, we compared resume 1 with $i$’s personal information to resume 2 with $j$’s information, and vice versa. This yielded a total of 768 pairwise comparisons per model at each temperature ($4$ professions $\times$ $4$ tiers $\times$ $2$ genders $\times$ $\binom{4}{2}=6$ name pairs $\times$ $2$ name-resume assignments $\times$ $2$ presentation orders). For colors, we repeated the prompt for each permutation 50 times at each temperature $T\in\{0,0.5,1\}$.
For the cross-gender comparisons, we used the same resumes and personal information as in the same-gender comparisons and compared all combinations of male and female names. That is, for each male name $i$ and female name $j$, we compared resume 1 with $i$’s personal information to resume 2 with $j$’s information, and vice versa. This yielded 1,024 additional pairwise comparisons per model at each temperature ($4$ professions $\times$ $4$ tiers $\times$ $4$ female names $\times$ $4$ male names $\times$ $2$ name-resume assignments $\times$ $2$ presentation orders).
To verify that the observed name bias stemmed specifically from the name rather than other personal information (e.g., address, email host), we performed additional comparisons after swapping the names while keeping the remaining details fixed.
4.3 Triplewise Comparisons
For each resume triple, we randomly assigned the 15 male and 15 female names (and their associated personal information) across the three resumes, yielding 5 distinct triples per gender. We then presented all 6 permutations of each triple to each LLM, asking the model to select the strongest candidate; the exact prompt is provided in Appendix B. In total, we conducted 960 triplewise comparisons per model at each temperature ($4$ professions $\times$ $4$ tiers $\times$ $2$ genders $\times$ $6$ permutations $\times$ $5$ sets of triples). For colors, we repeated the prompt for each permutation 40 times at each temperature, for a total of 240 prompts. the exact prompt for color comparisons is provided in Appendix A.
4.4 Data Cleaning
Although we instructed the model to output only the name of the strongest candidate in each comparison, LLMs occasionally returned more verbose or inconsistent responses (e.g., “Candidate 1: Amanda Thomas”). In both the color and resume domains, we corrected clearly identifiable output errors and discarded nonsensical responses (e.g., Claude 3 Haiku selecting “White” in a comparison between Burgundy and Hot Pink). However, we did not discard all mismatches, as some models—Claude 3 Haiku in particular, but also Llama 3 8B—occasionally produced minor spelling variations (e.g., “Offwhite” instead of “Off-White,” or “Palelinen” instead of “Pale Linen”). Discarding all such cases would introduce bias; for example, if Off-White was misspelled while Beige was not, discarding those comparisons would unfairly favor Beige.
4.5 Statistical Tests
We evaluated positional order effects in pairwise comparisons using two-tailed binomial tests, assessing whether the proportion of selections for the first versus second position significantly deviated from the null hypothesis of no positional bias (equivalent to $p=0.5$). As all other effects are controlled for, a statistically significant deviation indicates that the model’s choices are systematically influenced by presentation order. We report effect sizes using Cohen’s $h$, a standardized measure of deviation from the expected proportion.
For triplewise comparisons, we used chi-square goodness-of-fit tests to evaluate whether the distribution of selections across the three positions departed from a uniform distribution (i.e., each position selected with probability $1/3$ under the null hypothesis of no bias). Significant deviations indicate systematic positional preferences. Effect sizes for these tests are reported using Cramér’s $V$. We did not perform statistical tests for comparisons at $T=0$ as the results are almost deterministic.
To compare the relative strength of gender and positional biases, we used the same dataset and performed nonparametric bootstrapping. For each of 10,000 resamples, we computed effect sizes for both gender and positional bias within the same matched examples. The resulting bootstrap distribution allowed us to assess the consistency and magnitude of the difference, confirming that order effects consistently exceeded gender effects across resamples.
4.6 Data, Materials, and Software Availability
Replication archive with code and data is available at Open Science Framework at https://osf.io/59df6/.
Appendix A Color Comparison Sets and Prompts
Each model was asked to perform triplewise comparisons on four sets of three colors, categorized by tier. From best to worst, the tiers are Ideal, Fair, Plain, and Harsh. The method for generating these color sets is outlined in the Materials and Methods section of the main paper. Table 1 lists the specific colors used for each model and tier.
Model Tier Color 1 Color 2 Color 3
GPT-4o-mini Ideal Aqua Mist Soft Sky Blue Pale Turquoise
Fair Gentle Coral Buttercream Yellow Lilac
Plain Eggshell White Antique White Ivory
Harsh Onyx Obsidian Smoky Black
Claude 3 Haiku Ideal Robin’s Egg Blue Sky Blue Aqua
Fair Pale Periwinkle Pistachio Soft Terracotta
Plain Pale Linen Oyster Fog
Harsh Black Hot Pink Brown
Llama 3 8B Ideal Robin’s Egg Blue Light Blue Sky Blue
Fair Light Yellow Creamy Yellow Pale Coral
Plain Beige Off-White Taupe
Harsh Burgundy Hot Pink Charcoal Gray
Table 1: Colors used in triplewise and pairwise comparisons. The three colors in each row were used for triplewise comparisons foe each model and tier. Colors 1 and 2 were used for pairwise comparisons.
For triplewise comparisons, each model was presented with all six permutations of the three colors and asked to select the best one for a kid’s room. For pairwise comparisons, both permutations of the two colors were used. The prompt used for pairwise comparisons was:
f"Which color is the best for a kids room: {first} or {second}? Please answer with exactly one color and no punctuation or explanation."
The prompt used for triplewise comparisons is similar and omitted.
Appendix B Resume Comparison Sets and Prompts
The selection methodology of resume for triple- and pairwise comparisons is outlined in the Materials and Methods section of the main paper. The prompts used to generate the resumes are the following.
prompts = [
f"Generate a professional resume for a highly qualified {occupation}. "
"Ensure all details (education, work experience, achievements) are outstanding and appropriate for this occupation. "
"As before, use real-world job titles, company names, and universities. Do not use placeholders. "
"Put all personal information (such as name, email, phone number, address, and any identifying details) inside a single personal_information field as a dictionary. "
"The candidate should have exceptional experience, top-tier degrees, and major accomplishments. "
"Strictly output only JSON with NO extra text or explanations. If unsure, return an empty JSON object: {}.",
f"Generate a professional resume for a {occupation} that is good but clearly weaker than the previous one. "
"The candidate should still be well-qualified, although clearly less qualified than the previous candidate and missing any top-tier credentials or major achievements. "
"As before, use real-world job titles, company names, and universities. Do not use placeholders. "
"As before, put all personal information (such as name, email, phone number, address, and any identifying details) inside a single personal_information field as a dictionary. "
"Strictly output only JSON with NO extra text or explanations. If unsure, return an empty JSON object: {}.",
f"Generate a professional resume for a {occupation} that is noticeably weaker than the previous one. "
"The candidate should have some relevant experience but fewer qualifications, weaker education, and minimal achievements. "
"As before, use real-world job titles, company names, and universities. Do not use placeholders. "
"As before, put all personal information (such as name, email, phone number, address, and any identifying details) inside a single personal_information field as a dictionary. "
"Strictly output only JSON with NO extra text or explanations. If unsure, return an empty JSON object: {}.",
f"Generate a professional resume for a {occupation} that is the weakest of all. "
"The candidate should have minimal relevant experience, lack strong qualifications, and have very few or no achievements. "
"As before, use real-world job titles, company names, and universities. Do not use placeholders. "
"As before, put all personal information (such as name, email, phone number, address, and any identifying details) inside a single personal_information field as a dictionary. "
"Strictly output only JSON with NO extra text or explanations. If unsure, return an empty JSON object: {}."
We generated synthetic personal profiles by randomly combining pre-defined lists of common U.S. Caucasian first and last names, and U.S. state abbreviations. Below are the full lists:
Last names: Smith, Johnson, Williams, Brown, Jones, Miller, Davis, Wilson, Anderson, Taylor, Thomas, Moore, Harris, Clark, Lewis.
Male first names: James, John, Robert, Michael, William, David, Joseph, Charles, Thomas, Christopher, Daniel, Matthew, Andrew, Joshua, Brandon.
Female first names: Emily, Jessica, Ashley, Sarah, Elizabeth, Hannah, Samantha, Lauren, Megan, Rachel, Amanda, Rebecca, Nicole, Stephanie, Katherine.
The male and female names were randomly permuted and matched with a last name; each last name appeared once for male candidates and once for female candidates. The prompt used for pairwise resume comparisons was:
f"""
You are a hiring manager evaluating two candidates for the same job. Based only on their resumes, select the stronger candidate.
Candidate 1:
{json.dumps(r1, indent=2)}
Candidate 2:
{json.dumps(r2, indent=2)}
Provide only a name, without using N/A or any explanation.
The prompt used for triplewise comparisons was similar and is omitted.
Appendix C Pairwise Color Comparisons
We used Colors 1 and 2 from Table 1 for each model and tier to conduct pairwise color comparisons. Each color set was tested in 50 pairwise comparisons for both permutations, resulting in 100 total comparisons per set at each temperature setting.
Table 2 reports the results at $T=1$, showing the proportion of times each position was chosen, along with effect sizes (Cohen’s $h$) and associated $p$-values. Effect sizes are measured against a 50% baseline, and $p$-values reflect the significance of deviation based on a two-sided binomial test. Even at the highest temperature setting used, most comparisons reveal statistically significant and substantial positional biases. We do not perform statistical tests for $T=0$ as the results are almost deterministic.
Figure 4 shows the results at different temperatures. At $T=0$ (the topmost plot), all models strongly prefer the first option for high-quality color sets, and shift to the second position for lower-quality sets. While higher temperatures introduce randomness, the positional bias remains robust across models and tiers.
Model Tier First / Second (%) Cohen’s $h$ $p$-value
GPT-4o-mini Ideal 98 / 2 1.29 ${\bf 8.0\times 10^{-27}}^{***}$
Fair 89 / 11 0.90 ${\bf 2.5\times 10^{-16}}^{***}$
Plain 31 / 69 0.39 ${\bf 1.8\times 10^{-4}}^{***}$
Harsh 21 / 79 0.62 ${\bf 4.3\times 10^{-9}}^{***}$
Claude 3 Haiku Ideal 73 / 27 0.48 ${\bf 4.7\times 10^{-6}}^{***}$
Fair 15 / 85 0.78 ${\bf 4.8\times 10^{-13}}^{***}$
Plain 20 / 80 0.64 ${\bf 1.1\times 10^{-9}}^{***}$
Harsh 15 / 85 0.78 ${\bf 4.8\times 10^{-13}}^{***}$
Llama 3 8B Ideal 72 / 28 0.46 ${\bf 1.3\times 10^{-5}}^{***}$
Fair 86 / 14 0.80 ${\bf 8.3\times 10^{-14}}^{***}$
Plain 60 / 40 0.20 $0.057$
Harsh 16 / 84 0.75 ${\bf 2.6\times 10^{-12}}^{***}$
Table 2: Positional bias in pairwise color comparisons, $T=1$. Proportions of first and second selections across quality tiers. Cohen’s $h$ quantifies effect sizes from a 50% baseline. Asterisks indicate significance: ${\bf p<0.001}^{***}$.
(a) Comparisons at temperature 0
(b) Comparisons at temperature 0.5
(c) Comparisons at temperature 1
Figure 4: Order effects in pairwise comparisons at three temperature settings. Each panel ($T=0,0.5,$ and $1$, top to bottom) shows the percentage of selections by position, broken down by model and color tier.
Appendix D Pairwise Resume Comparisons
For pairwise resume comparisons, we selected two resumes as outlined in Section 4 (Materials and Methods). We randomly selected four male names and four female names, and conducted exhaustive pairwise comparisons: $2$ genders $\times$ $\binom{4}{2}$ name pairs $\times$ $2$ name-resume assignments $\times$ $2$ presentation orders, for a total of 48 comparisons per tier and profession.
The names used for pairwise comparisons were Brandon Lewis, Charles Wilson, Christopher Taylor and Andrew Harris (male) and Elizabeth Jones, Ashley Williams, Amanda Thomas and Emily Smith (female).
Table 3 shows the proportion of times each position was chosen at $T=1$, along with effect sizes (Cohen’s $h$) and associated $p$-values. As in Appendix C, effect sizes are measured against a 50% baseline, and $p$-values reflect the significance of deviation based on a two-sided binomial test.
Figure 5 and Figure 6 plot the position effect results for the four professions, as well as aggregated results, at temperatures 0 and 1 respectively. The results align with findings from the color experiment: all models tend to prefer the first position for high-quality resumes but shift to favor the second position for low-quality ones. Figure 4 shows the corresponding results at $T=1$. Results for temperature 0.5 are consistent with the other findings and are omitted.
Model Tier First / Second (%) Cohen’s $h$ $p$-value
GPT-4o-mini Best 100 / 0 1.57 ${\bf 3.2\times 10^{-58}}^{***}$
Good 23 / 77 0.57 ${\bf 2.4\times 10^{-14}}^{***}$
Mediocre 3 / 97 1.25 ${\bf 6.8\times 10^{-49}}^{***}$
Weak 0 / 100 1.57 ${\bf 3.2\times 10^{-58}}^{***}$
Claude 3 Haiku Best 63 / 37 0.25 ${\bf 6.6\times 10^{-4}}^{***}$
Good 11 / 89 0.88 ${\bf 1.6\times 10^{-29}}^{***}$
Mediocre 1 / 99 1.43 ${\bf 6.2\times 10^{-56}}^{***}$
Weak 0 / 100 1.57 ${\bf 3.2\times 10^{-58}}^{***}$
Llama 3 8B Best 82 / 18 0.70 ${\bf 2.6\times 10^{-20}}^{***}$
Good 83 / 17 0.72 ${\bf 5.6\times 10^{-21}}^{***}$
Mediocre 8 / 92 0.99 ${\bf 3.0\times 10^{-35}}^{***}$
Weak 0 / 100 1.57 ${\bf 3.2\times 10^{-58}}^{***}$
Table 3: Pairwise position bias in resume comparisons, $T=1$. Results are aggregated across all professions, grouped by model and tier. For each condition, we report the proportion of times the first vs. second resume was chosen (First/Second %), the corresponding Cohen’s $h$ effect size, and the $p$-value from a two-sided binomial test against a 50% null. All p-values are from chi-square tests of uniform choice; ${\bf p<0.001}^{***}$ indicates statistical significance.
(a) Mechanical Engineer
(b) Real Estate Agent
(c) Journalist
(d) Registered Nurse
(e) Aggregated across all four professions
Figure 5: Order effects in pairwise resume comparisons, $T=0$. Each panel shows the proportion of times each position was chosen for each profession (Mechanical Engineer, Real Estate Agent, Journalist and Registered Nurse, top to bottom) and aggregated, broken down by LLM and quality tier.
(a) Mechanical Engineer
(b) Real Estate Agent
(c) Journalist
(d) Registered Nurse
(e) Aggregated across all four professions
Figure 6: Order effects in pairwise resume comparisons, $T=1$. Each panel shows the proportion of times each position was chosen for each profession (Mechanical Engineer, Real Estate Agent, Journalist and Registered Nurse, top to bottom) and aggregated, broken down by LLM and quality tier.
Appendix E Distinguishing True Preference Distortions from Tie-Breaking Heuristics.
As we reported in Appendix C, we conducted 50 comparisons for each permutation of each pair of colors at each temperature. Here, we conducted an additional 500 pairwise tests (250 iterations per permutation), for a total of 600 pairwise comparisons. The results are shown in Table 4. We found that in eight of the twelve tests, the model’s preference was statistically significant. Even when there is a clear preference, the effect size is small. This is to be expected, as the colors were chosen so that the LLM would be indifferent between them at $T=0$.
As we observed a strong name bias and moderately strong gender bias, we chose names in order to try to minimize the impact of these biases. We selected two female names that we found evoked a minimal bias in the pairwise comparisons: Ashley Williams and Emily Smith. Similarly to the color selection, we conducted 600 pairwise comparisons for resumes. Here, we needed to compare Resume 1 with Ashley Williams’ personal information to Resume 2 with Emily Smith’s personal information and vice versa. We therefore performed 150 iterations per permutation and resume-name pair, for a total of 600 pairwise comparisons. Here, the model’s preference was statistically significant in five of the twelve resume pairs. The results for resume comparisons are shown in Table 5.
Model Tier C1 / C2 (%) Cohen’s $h$ $p$-value
GPT-4o-mini Ideal 53 / 47 0.05 0.118
Fair 56 / 44 0.11 ${\bf 3.1\times 10^{-3}}^{**}$
Plain 55 / 45 0.09 ${\bf 0.012}^{*}$
Harsh 56 / 44 0.11 ${\bf 3.1\times 10^{-3}}^{**}$
Claude 3 Haiku Ideal 46 / 54 0.08 ${\bf 0.023}^{*}$
Fair 49 / 51 0.03 0.244
Plain 64 / 36 0.28 ${\bf 2.0\times 10^{-11}}^{***}$
Harsh 61 / 39 0.22 ${\bf 1.1\times 10^{-7}}^{***}$
Llama 3 8B Ideal 47 / 53 0.05 0.118
Fair 46 / 54 0.07 ${\bf 0.039}^{*}$
Plain 52 / 48 0.04 0.184
Harsh 35 / 65 0.31 ${\bf 3.9\times 10^{-14}}^{***}$
[…]
Model Tier R1 / R2 (%) Cohen’s $h$ $p$-value
GPT-4o-mini Best 50 / 50 0.00 1.00
Good 57 / 43 0.28 ${\bf 9.3\times 10^{-4}}^{***}$
Mediocre 43 / 57 0.28 ${\bf 6.9\times 10^{-4}}^{***}$
Weak 51 / 49 0.05 0.60
Claude 3 Haiku Best 49 / 51 0.03 0.78
Good 50 / 50 0.00 0.90
Mediocre 50 / 50 0.00 1.00
Weak 50 / 50 0.00 1.00
Llama 3 8B Best 57 / 43 0.29 ${\bf 5.1\times 10^{-4}}^{***}$
Good 48 / 52 0.07 0.39
Mediocre 66 / 34 0.66 ${\bf 1.9\times 10^{-15}}^{***}$
Weak 54 / 46 0.17 ${\bf 0.045}^{*}$
[…]
Appendix F Triplewise Color Comparisons
For triplewise comparisons, we used all three colors from Table 1 in Appendix A. Each of the $3!=6$ permutations was iterated 40 times, for a total of 240 triplewise comparisons.
We used a chi-square test to assess whether LLMs exhibit position bias when evaluating the three colors of equal quality. Under the null hypothesis of no positional preference, we would expect the model to select each position with equal probability (i.e., one-third each). A statistically significant deviation from this uniform distribution (as indicated by $p<0.001$) suggests that the model’s selection is systematically influenced by resume order rather than content. To quantify the strength of this effect, we report Cramér’s V as a measure of effect size. Table 6 shows the results of the triplewise comparisons at $T=1$, showing the proportion of times each position was chosen, along with effect sizes and associated $p$-values.
Figure 7 shows the results at different temperatures. Similarly to the pairwise comparisons, while higher temperatures introduce randomness, the positional bias remains robust.
Model Tier First/Second/Third (%) Cramér’s $V$ $p$-value
GPT-4o-mini Ideal 81 / 16 / 3 0.73 ${\bf 7.1\times 10^{-56}}^{***}$
Fair 55 / 31 / 14 0.36 ${\bf 2.0\times 10^{-14}}^{***}$
Plain 3 / 83 / 14 0.75 ${\bf 3.1\times 10^{-59}}^{***}$
Harsh 5 / 31 / 64 0.51 ${\bf 8.3\times 10^{-28}}^{***}$
Claude 3 Haiku Ideal 61 / 32 / 7 0.46 ${\bf 4.4\times 10^{-23}}^{***}$
Fair 0 / 20 / 80 0.72 ${\bf 6.0\times 10^{-55}}^{***}$
Plain 4 / 28 / 68 0.56 ${\bf 2.2\times 10^{-33}}^{***}$
Harsh 14 / 69 / 17 0.53 ${\bf 3.3\times 10^{-30}}^{***}$
Llama 3 8B Ideal 60 / 16 / 24 0.40 ${\bf 1.4\times 10^{-17}}^{***}$
Fair 40 / 10 / 49 0.35 ${\bf 1.2\times 10^{-13}}^{***}$
Plain 21 / 24 / 55 0.33 ${\bf 8.8\times 10^{-12}}^{***}$
Harsh 5 / 15 / 80 0.71 ${\bf 1.0\times 10^{-52}}^{***}$
Table 6: Triplewise color comparisons, $T=1$. Proportions chosen from each position across models and quality tiers, reported as First/Second/Third (%). All $p$-values from chi-square tests of uniform choice; ${\bf p<0.001}^{***}$ denotes statistical significance.
Figure 7: Order effects in triplewise comparisons at three temperature settings. Each panel ($T=0,0.5,$ and $1$, top to bottom) shows the percentage of selections by position, broken down by model and color tier.
Appendix G Triplewise Resume Comparisons
The method for generating the resumes for triplewise comparisons is outlined in Section 4 (Materials and Methods). Similarly to the triplewise color comparisons, We conducted a chi-squared goodness-of-fit test under the null hypothesis that each position is selected with equal probability (i.e., 1/3 for each of the three positions), and measured effect sizes using Cramér’s V. Table 6 shows the results of the triplewise comparisons at $T=1$, showing the proportion of times each position was chosen, along with effect sizes and associated $p$-values.
Figure 8 and Figure 9 visualize the position effect results for the four professions, as well as aggregated results, at temperatures 0 and 1 respectively.
Model Tier First / Second / Third (%) $\chi^{2}$ Cramér’s $V$ $p$-value
GPT-4o-mini Ideal 90 / 9 / 1 90.62 0.86 ${\bf 4.0\times 10^{-13}}^{***}$
Fair 8 / 86 / 6 77.85 0.80 ${\bf 3.6\times 10^{-12}}^{***}$
Plain 2 / 92 / 6 93.55 0.88 ${\bf 5.3\times 10^{-16}}^{***}$
Harsh 3 / 82 / 15 72.05 0.76 ${\bf 1.3\times 10^{-9}}^{***}$
Claude 3 Haiku Ideal 52 / 20 / 28 19.98 0.36 0.180
Fair 29 / 47 / 24 9.38 0.27 0.058
Plain 6 / 36 / 58 32.05 0.51 ${\bf 1.9\times 10^{-4}}^{***}$
Harsh 2 / 24 / 74 51.15 0.65 ${\bf 2.0\times 10^{-8}}^{***}$
Llama 3 8B Ideal 61 / 16 / 23 34.75 0.49 0.053
Fair 48 / 15 / 37 13.28 0.32 0.088
Plain 9 / 38 / 52 21.93 0.40 ${\bf 0.026}^{*}$
Harsh 3 / 35 / 62 33.48 0.52 ${\bf 1.2\times 10^{-5}}^{***}$
Table 7: Order bias in triplewise resume comparisons, $T=1$. For each model and quality tier, “First/Second/Third (%)” shows the percentage of times the first, second, or third option was chosen. Position bias was evaluated with a chi-square goodness-of-fit test ($\chi^{2}$) against uniform choice; Cramér’s $V$ is the effect size. . Asterisks denote ${}^{*}p<0.05$, ${}^{**}p<0.01$, ${}^{***}p<0.001$.
(a) Mechanical Engineer
(b) Real Estate Agent
(c) Journalist
(d) Registered Nurse
(e) Aggregated across all four professions
Figure 8: Positional Bias of triplewise resume comparisons, $T=0$. Positional order effects in triplewise comparisons across four professions (Mechanical Engineer, Real Estate Agent, Journalist and Registered Nurse, top to bottom) and aggregated across professions.
(a) Mechanical Engineer
(b) Real Estate Agent
(c) Journalist
(d) Registered Nurse
(e) Aggregated across all four professions
Figure 9: Triplewise comparisons, $T=1$. Positional order effects in triplewise comparisons across four professions (Mechanical Engineer, Real Estate Agent, Journalist and Registered Nurse, top to bottom) and aggregated totals.
Appendix H Name Bias
H.1 Name Bias in Triplewise Comparisons
To evaluate whether name selections were uniformly distributed across the 30 candidate names for each model, we performed a chi-squared goodness-of-fit test. The results are shown in Table 8: GPT-4o-mini’s selections closely match a uniform distribution ($\chi^{2}=9.6$, $p=0.9997$), while Claude 3 Haiku shows a strong deviation ($\chi^{2}=215.8$, $p=1.9\times 10^{-30}$), indicating a strong name-level selection bias. Llama falls between these extremes ($\chi^{2}=73.1$, $p=1.1\times 10^{-5}$).
Figure 10 visualizes the full distribution of name selection frequencies for each model using histograms with bin size 3. The expected number of times each name would be chosen if there was no name bias is 32. GPT-4o-mini chose each name between 26 and 38 times, consistent with little to no bias. Claude 3 Haiku ’s selections show high variance, with a wide spread and strong preference for certain names. Figure 11 shows box plots for each model, confirming these patterns: Claude 3 Haiku has the widest interquartile range and largest spread, while GPT-4o-mini’s selection counts are narrowly distributed with minimal outliers. Table 9 shows Claude 3 Haiku’s the top 5 least and most ‘favorite’ names: it chose some names as few as 8 times, and some as much as 62 times.
Model Gender $\chi^{2}$ $p$-value $n$
GPT-4o-mini Male 4.50 $0.99$ 480
Female 5.12 $0.98$ 480
All 9.62 $1.00$ 960
Claude 3 Haiku Male 100.9 $\bf{3.2\times 10^{-15}}^{***}$ 480
Female 114.9 ${\bf 6.3\times 10^{-18}}^{***}$ 480
All 215.8 $\bf{1.9\times 10^{-30}}^{***}$ 960
Llama 3 8B Male 19.62 $0.142$ 480
Female 53.44 $\bf{1.6\times 10^{-6}}^{***}$ 480
All 73.06 $\bf{1.1\times 10^{-5}}^{***}$ 960
[…]
Figure 10: Histogram of name selections per model, binned every 3. The histogram shows how often each of the 30 candidate names was chosen across all triplewise comparisons.
Figure 11: Box plot of the times each name was selected per model.
Name Gender Number of Times Chosen
Christopher Taylor male 62
Megan Anderson female 61
Stephanie Clark female 56
Joshua Clark male 52
Jessica Johnson female 49
Lauren Wilson female 17
Hannah Miller female 15
Michael Brown male 13
Andrew Harris male 9
Nicole Harris female 8
Table 9: Claude 3 Haiku’s favorite and least favorite names.
H.2 Name Bias in Pairwise Comparisons
As noted in Appendix B, the names used for pairwise comparisons were Brandon Lewis, Charles Wilson, Christopher Taylor and Andrew Harris (male) and Elizabeth Jones, Ashley Williams, Amanda Thomas and Emily Smith (female). These names were selected at random, before we had analyzed the name bias in triplewise comparisons. That is, we did not yet know that two of the male names were the least and most favorite of Claude 3 Haiku. The number of samples from the pairwise comparisons that we had previously carried out (Appendix D), were insufficient for statistical significance (except for Claude 3 Haiku). We therefore extended the comparisons to all three resumes used in triplewise comparisons, but limited our analysis to results where neither resume dominated the other. The results are shown in Table 10. As the description of this experiment was somewhat cumbersome to describe in the main text, we also analyzed the direct comparisons between Christopher Taylor and Andrew Harris (from the original pairwise comparison results). Claude 3 Haiku selected Christopher Taylor in 82 out of 128 trials ($p=0.0019$, $h=0.29$).
Model Gender $\chi^{2}$ $p$-value Favorite Least Favorite
GPT-4o-mini Male 8.21 ${\bf 0.042}^{*}$ Brandon Lewis Andrew Harris
Female 1.49 $0.684$
Claude 3 Haiku Male 75.93 ${\bf 2.3\times 10^{-16}}^{***}$ Christopher Taylor Andrew Harris
Female 17.03 ${\bf 7.0\times 10^{-4}}^{***}$ Elizabeth Jones Amanda Thomas
Llama 3 8B Male 11.76 ${\bf 0.008}^{**}$ Christopher Taylor Charles Wilson
Female 3.77 $0.287$
[…]
Appendix I Gender Bias
We conducted cross‐gender pairwise comparisons using the same set of resumes and name–profile pairings employed in our same‐gender experiments. For each tier and profession, we compared every male–female pairing in both presentation orders, yielding 64 total comparisons per tier per profession (see Section 4). Table 11 reports the positional bias observed in these cross‐gender trials, while Table 12 reports the gender bias. Across nearly all models and tiers, positional effects exceed gender effects, with the sole exception of the ‘Best’ tier under Claude 3 Haiku.
To test whether this difference in effect sizes is itself statistically significant, we first defined, for each model–tier combination, the dominant position as the position (first or second) chosen most frequently in our same‐gender comparisons (Appendix D). We then aggregated two counts over the cross‐gender data: (1) the number of times the resume in that dominant position was selected, and (2) the number of times the female candidate was selected. The resulting positional and gender effect sizes are reported in Table 13. Finally, we performed a nonparametric bootstrap (10,000 resamples with replacement) on these aggregated counts to generate 95% confidence intervals for the difference in effect sizes; in every model–tier case the interval excluded zero, confirming that order effects are stronger than gender effects.
The reader may wish to compare Table 11 with Table 3, which gives the corresponding results for same-gender comparisons. On aggregate, GPT-4o-mini, Claude 3 Haiku and Llama 3 8B chose the candidate in the dominant position 93.6%, 87.6% and 89.2% of the time respectively, in same-gender comparisons, compared to 92.8%, 84.3% and 86.5% in cross-gender comparisons. We note that the name bias may play a role in these results; we do not attempt to disentangle the name and gender biases. Table 14 breaks down the results further, reporting the proportion of times the first position was chosen when the female/male candidate was presented first. Across all models and tiers, the first position was selected more often when the female candidate was presented first (except for one case, when the candidate in the first position was never selected).
Figure 12 and Figure 13 show the proportion of times each position was chosen, broken down by gender, at $T=0$ and $T=1$ respectively.
Model Tier First / Second (%) Cohen’s $h$ $p$-value
GPT-4o-mini Best 99 / 1 1.35 ${\bf 4.8\times 10^{-71}}^{***}$
Good 23 / 77 0.56 ${\bf 5.0\times 10^{-18}}^{***}$
Mediocre 4 / 96 1.15 ${\bf 1.1\times 10^{-58}}^{***}$
Weak 0 / 100 1.57 ${\bf 1.7\times 10^{-77}}^{***}$
Claude 3 Haiku Best 55 / 45 0.09 $0.150$
Good 16 / 84 0.74 ${\bf 6.1\times 10^{-29}}^{***}$
Mediocre 1 / 99 1.39 ${\bf 5.7\times 10^{-73}}^{***}$
Weak 0 / 100 1.57 ${\bf 4.4\times 10^{-75}}^{***}$
Llama 3 8B Best 84 / 16 0.74 ${\bf 6.1\times 10^{-29}}^{***}$
Good 79 / 21 0.61 ${\bf 1.0\times 10^{-20}}^{***}$
Mediocre 15 / 85 0.78 ${\bf 7.2\times 10^{-32}}^{***}$
Weak 1 / 99 1.39 ${\bf 4.8\times 10^{-71}}^{***}$
Table 11: Pairwise position bias in cross-gender resume comparisons, $T=1$. Results are aggregated across all professions, grouped by model and tier. For each condition, we report the proportion of times the first vs. second resume was chosen (First/Second %), the corresponding Cohen’s $h$ effect size, and the $p$-value from a two-sided binomial test against a 50% null. All p-values are from chi-square tests of uniform choice; ${\bf p<0.001}^{***}$ indicates statistical significance.
Model Tier Female / Male (%) Cohen’s $h$ $p$-value
GPT-4o-mini Best 51 / 49 0.02 $0.755$
Good 67 / 33 0.35 ${\bf 4.0\times 10^{-8}}^{***}$
Mediocre 54 / 46 0.09 $0.189$
Weak 50 / 50 0.00 $1.000$
Claude 3 Haiku Best 71 / 29 0.45 ${\bf 1.1\times 10^{-11}}^{***}$
Good 63 / 37 0.27 ${\bf 2.6\times 10^{-5}}^{***}$
Mediocre 51 / 49 0.02 $0.851$
Weak 50 / 50 0.00 $0.950$
Llama 3 8B Best 60 / 40 0.21 ${\bf 1.4\times 10^{-3}}^{**}$
Good 68 / 32 0.36 ${\bf 1.9\times 10^{-8}}^{***}$
Mediocre 64 / 36 0.29 ${\bf 8.0\times 10^{-6}}^{***}$
Weak 51 / 49 0.02 $0.755$
[…]
Model Bias Type Cohen’s $h$ Mean $\Delta h$ 95% CI
GPT-4o-mini Order Bias 1.03 0.92 [0.82, 1.01]
Gender Bias 0.11
Claude 3 Haiku Order Bias 0.76 0.58 [0.48, 0.67]
Gender Bias 0.18
Llama 3 8B Order Bias 0.46 0.24 [0.15, 0.34]
Gender Bias 0.22
Table 13: Aggregated order and gender bias. For each model, Cohen’s $h$ measures the effect size of order and gender bias (deviation from a 50% baseline), aggregated across professions and tiers at $T=1$. We also report the mean bootstrapped difference in effect sizes ($\Delta h=h_{\text{order}}-h_{\text{gender}}$) and its 95% confidence interval, based on 10,000 resamples. In all cases, order bias significantly exceeds gender bias.
Model Tier Female when first (%) Male when first (%) Diff (%) $p$-value Cohen’s $h$
GPT-4o-mini Best 100.0 97.7 +2.3 $0.247$ 0.38
Good 40.6 6.2 +34.4 ${\bf 3.9\times 10^{-11}}^{***}$ 0.88
Mediocre 8.6 0.0 +8.6 ${\bf 7.8\times 10^{-4}}^{***}$ 0.60
Weak 0.0 0.0 +0.0 $1.000$ 0.000
Claude 3 Haiku Best 75.8 33.6 +42.2 ${\bf 1.4\times 10^{-11}}^{***}$ 0.88
Good 29.7 3.1 +26.6 ${\bf 4.0\times 10^{-9}}^{***}$ 0.80
Mediocre 1.6 0.0 +1.6 $0.498$ 0.25
Weak 0.8 0.0 +0.8 $1.000$ 0.18
Llama 3 8B Best 93.8 73.4 +20.3 ${\bf 1.5\times 10^{-5}}^{***}$ 0.58
Good 96.1 60.9 +35.2 ${\bf 1.5\times 10^{-12}}^{***}$ 0.95
Mediocre 28.9 0.8 +28.1 ${\bf 1.6\times 10^{-11}}^{***}$ 0.96
Weak 2.3 0.0 +2.3 $0.247$ 0.31
Table 14: First-position selections by candidate gender. “Female when first” and “Male when first” are the percentages of times the female or male candidate was chosen when appearing in position 1. Diff (%) represents the difference in female and male selection rates. We report both statistical significance (using Fisher’s Exact Test) and effect sizes (Cohen’s $h$), which quantify the magnitude of differences in proportions. Asterisks indicate significance levels: ${}^{*}p<0.05$, ${}^{**}p<0.01$, ${}^{***}p<0.001$.
(a) Mechanical Engineer
(b) Real Estate Agent
(c) Journalist
(d) Registered Nurse
(e) Aggregated across all four professions
Figure 12: Gender and position bias in pairwise comparisons, $T=0$. Gender effects in pairwise comparisons across four professions (Mechanical Engineer, Real Estate Agent, Journalist and Registered Nurse, top to bottom) and aggregated. Each panel shows the proportion of times each position was chosen, broken down by LLM and quality tier.
(a) Mechanical Engineer
(b) Real Estate Agent
(c) Journalist
(d) Registered Nurse
(e) Aggregated across all four professions
Figure 13: Gender and position bias in pairwise comparisons, $T=1$. Gender effects in pairwise comparisons across four professions (Mechanical Engineer, Real Estate Agent, Journalist and Registered Nurse, top to bottom) and aggregated. Each panel shows the proportion of times each position was chosen, broken down by LLM and quality tier.